sqrt(9x+40)-6=2x-1
I have no idea what to do. I have tried multiple things and nothing gives me 1 which is the answer my professor said is the correct one.Help.Thank You

Answer by scott8148(6628)

adding 6 ___ sqrt(9x+40) =2x + 5

squaring ___ 9x + 40 = 4x^2 + 20x + 25

subtracting (9x + 40) ___ 0 = 4x^2 + 11x - 15

factoring ___ 0 = (4x + 15)(x - 1)

4x + 15 = 0 ___ x = -15/4

x - 1 = 0 ___ x = 1